OpenStudy (anonymous):

Estimate the derivative of f(x0=2^x at t=3

OpenStudy (anonymous):

i suppose your question is f(x)=2^x at x=3 well f(3)=2^3=8

OpenStudy (anonymous):

but I have a question cause I thought that to but the question in my book actually says t=3 can I change that to x=3

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

i think it's type error

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Ok now I am second quessing my self cause the question also says that the answers value should be accurate to two decimal places

OpenStudy (anonymous):

are you sure you put in all the questions right?

OpenStudy (anonymous):

yes it is Estimate the derivative of f(x)=2^x at t=3 the answers value should be accurate to two decimal places. that is the question

OpenStudy (anonymous):

You know the answer is going to be \(8\ln(2)\)... So is it just a matter of calculating \(\ln(2)\)?

OpenStudy (anonymous):

so then are you saying the answer is 5.55

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

wio is that what you got

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

thanks